### v1

Abstract
Hyperparameter optimization (HPO) plays a central role in the performance of deep learning models, yet remains computationally expensive and difficult to interpret, particularly for time-series forecasting. While Bayesian Optimization (BO) is a standard approach, it typically treats tuning tasks independently and provides limited insight into its decisions. Recent advances in large language models (LLMs) offer new opportunities to incorporate structured prior knowledge and reasoning into optimization pipelines. We introduce LLM-AutoOpt, a hybrid HPO framework that combines BO with LLM-based contextual reasoning. The framework encodes dataset meta-features, model descriptions, historical optimization outcomes, and target objectives as structured meta-knowledge within LLM prompts, using BO to initialize the search and mitigate cold-start effects. This design enables context-aware and stable hyperparameter refinement while exposing the reasoning behind optimization decisions. Experiments on a multivariate time series forecasting benchmark demonstrate that LLM-AutoOpt achieves improved predictive performance and more interpretable optimization behavior compared to BO and LLM baselines without meta-knowledge.
1 Introduction
AutoML is the process of automating the tasks involved in building and deploying machine learning models [27]. One of the most challenging components of AutoML is hyperparameter optimization (HPO), which is essential for developing robust and effective deep learning models. The goal of HPO is to identify optimal configurations that enhance model performance, stability, and interpretability. Traditional HPO methods such as grid search, random search, Bayesian optimization, and gradient-based approaches are commonly used to optimize hyperparameters and have demonstrated improvements in model accuracy [28]. However, these methods still struggle with large-scale data and complex models including deep neural networks, and are often computationally expensive [1, 2, 6]. AutoML enhances the hyperparameter optimization task by incorporating different learning paradigms such as meta-learning and reinforcement learning.
Meta-learning enables hyperparameter optimization (HPO) methods to rapidly adapt to new tasks by leveraging experience from previously observed tasks, where each task is represented through informative meta-features [3]. By transferring prior optimization knowledge, meta-learning–based HPO significantly improves both sample efficiency and adaptability compared to conventional approaches that treat each task independently. Crucially, these methods mitigate the need to initialize the optimization process from scratch, instead warm-starting the search using insights derived from historical tasks [4, 8, 9, 10]. This paradigm has been successfully operationalized in several AutoML systems, most notably Auto-sklearn [5], demonstrating its practical effectiveness in real-world settings.
As large language models (LLMs) have become increasingly integrated into machine learning research, they have also been applied to hyperparameter optimization (HPO). Prior work has leveraged LLMs to automate tasks such as neural architecture search and end-to-end ML pipeline construction [11, 12]. Owing to their strong in-context learning capabilities, LLMs can rapidly adapt to new tasks without explicit fine-tuning [13, 14]. This property has recently been extended to optimization settings, where LLMs are explored as alternatives or complements to traditional HPO methods. For example, Bayesian optimization has been reformulated as a natural language reasoning task, enabling LLMs to exploit their contextual reasoning abilities to recommend improved hyperparameter configurations [17]. More recent approaches, such as MetaLLMix [23], integrate meta-learning with LLM reasoning and explainable AI to transfer knowledge across tasks using historical experiment meta-datasets. In this framework, LLMs reason over structured meta-information to generate task-specific model and hyperparameter recommendations, with SHAP-based explanations enhancing transparency. While such methods have shown promising results for classification and regression tasks [22], their applicability to time-series forecasting remains largely unexplored due to the presence of strong temporal dependencies and evolving data distributions. Prior work has operationalized meta-learning for hyperparameter optimization through several complementary strategies. Feurer et al. [9] demonstrate that Bayesian optimization can be effectively warm-started by conditioning surrogate models on dataset meta-features, thereby guiding the search toward promising regions of the hyperparameter space. In parallel, In [8], the authors propose transferring high-performing configurations from meta-datasets of previous tuning runs to initialize optimization on new tasks. Extending these ideas, Lindauer and Hutter [10] show that incorporating historical benchmark data into warm-started, model-based configuration frameworks improves scalability while preserving optimization efficiency. Collectively, these methods highlight how meta-knowledge can be systematically exploited to accelerate convergence and improve solution quality relative to cold-start optimization.
To address this, we propose LLM-AutoMOpt, a framework for automating hyperparameter optimization in deep learning models for time series forecasting. The framework follows an iterative optimization process in which Bayesian optimization is first employed to initialize the search space and provide informative prior configurations, which are subsequently refined through structured LLM-guided reasoning. The LLM operates on carefully constructed prompts enriched with meta-knowledge, including dataset characteristics, model descriptions, target loss objectives, and historical optimization outcomes. This structured meta-information enables the model to generate context-aware and performance-driven hyperparameter recommendations. A key challenge lies in prompt design, which must balance expressive reasoning with well-defined constraints to mitigate hallucination and ensure stable optimization behavior. Beyond identifying optimal configurations, LLM-AutoMOpt emphasizes transparency and interpretability, explicitly exposing the meta-knowledge and reasoning processes underlying its optimization decisions.
2 Related works
Hyperparameter optimization (HPO) seeks to automatically identify hyperparameter configurations that maximize a model’s performance on a given task, often at the cost of expensive and time-consuming model evaluations [27]. Early approaches such as grid search systematically enumerate predefined configurations, guaranteeing coverage of the search space but becoming computationally prohibitive as the number or range of hyperparameters grows [36]. Random search addresses this limitation by sampling configurations independently, frequently achieving superior efficiency—particularly when only a small subset of hyperparameters has a dominant effect on performance [37]. Bayesian optimization further advances HPO by learning a probabilistic surrogate of the objective function and selecting new configurations via an explicit exploration–exploitation trade-off, enabling faster convergence with significantly fewer evaluations [29]. Despite their success, these methods generally treat each optimization problem independently, failing to exploit knowledge from prior tasks [1]. This limitation reduces scalability and sample efficiency, especially in computationally demanding settings such as time-series forecasting, where training costs are high and model behavior is strongly data-dependent [2].
To overcome this limitation, meta-learning has emerged as an effective paradigm for improving scalability and resource efficiency across a wide range of learning problems [30]. By leveraging knowledge acquired from previously solved tasks, meta-learning enables hyperparameter optimization to adapt more rapidly to new tasks, effectively avoiding repeated optimization from scratch. In particular, Bayesian optimization has been shown to benefit significantly from meta-learning–driven warm starts, where prior task information guides the initial search process [31, 32]. This combination has been successfully applied to complex attention-based architectures, leading to improved forecasting accuracy in energy systems compared to standalone optimization approaches [33]. Building on this idea, meta-Bayesian optimization learns informative surrogate model priors from related tasks, further accelerating convergence and improving optimization efficiency on new problems [38].
Beyond Bayesian and meta-learning–based approaches, reinforcement learning (RL) has also been explored for hyperparameter optimization by formulating HPO as a sequential decision-making problem. In this setting, an RL agent iteratively selects hyperparameter configurations based on feedback from previous evaluations, learning a policy that aims to maximize expected performance while reducing the number of costly evaluations [34]. Building on this idea, meta-learning has been combined with RL to further improve efficiency. For example, MultiTaskPB2 leverages metadata from prior RL training runs to warm-start hyperparameter optimization and bias surrogate models for new tasks, achieving significantly faster and more effective optimization than standard methods [35].
More recently, large language models (LLMs) have introduced a new paradigm for hyperparameter optimization by leveraging natural language understanding, contextual reasoning, and few-shot generalization capabilities [15, 16, 17, 18, 19]. In AgentHPO [15], an LLM-based agent autonomously analyzes task descriptions and historical trials, proposes new hyperparameter configurations, executes experiments, and iteratively refines its recommendations, forming a human-like and interpretable optimization loop that reduces manual intervention. LLMs have also been integrated with traditional HPO strategies to improve efficiency. For instance, coarse grid search can be guided by an LLM to identify and refine promising regions of the search space, effectively pruning unlikely configurations and reducing computational cost while maintaining or improving tuning quality [18]. Furthermore, LLMs have been combined with Bayesian optimization in frameworks such as LLAMBO, where the BO process is reformulated as a natural language reasoning task. This enables LLMs to propose and evaluate candidate configurations conditioned on past trials, supporting zero-shot warm-starting, improved surrogate modeling, and more focused exploration of promising regions, ultimately leading to faster convergence and superior hyperparameter tuning compared to conventional BO [17]. Beyond directly steering the optimization process, recent research has investigated how LLMs can be used as meta-learners that reason over collections of prior experiments. Frameworks such as MetaLLMix [22] construct structured meta-databases from historical training runs and rely on LLMs to synthesize transferable optimization knowledge across tasks. Related in-context meta-learning approaches [23] encode datasets and experimental settings as structured, human-readable metadata, enabling LLMs to infer suitable hyperparameter configurations in zero-shot or few-shot regimes.
Unlike agent-based or BO-integrated LLM methods, these approaches emphasize knowledge transfer at the task level rather than iterative trial-by-trial optimization. However, they have primarily been evaluated on classification and regression benchmarks and do not explicitly address the domain-specific challenges of time-series forecasting, where strong temporal dependencies, heterogeneous sampling rates, seasonal dynamics, and high training costs require specialized meta-features and optimization strategies. To address these limitations, we propose LLM-AutoMOpt, a forecasting-specific HPO framework that tightly integrates Bayesian optimization for structured initialization with LLM-based in-context reasoning over rich temporal meta-knowledge. In contrast to prior work, LLM-AutoMOpt explicitly accounts for temporal structure and model dynamics, enabling more sample-efficient, interpretable, and adaptive hyperparameter optimization for deep sequence models such as LSTM, Bi-LSTM, and Transformer architectures.
3 Methodology
LLM-AutoOpt is a hyperparameter optimization framework that exploits the reasoning capabilities of large language models (LLMs) to recommend high-quality hyperparameter configurations for deep learning models in time series forecasting. In addition to improving optimization efficiency, the framework emphasizes transparency by making the optimization process interpretable. LLM-AutoOpt operates iteratively, progressively refining the LLM’s contextual understanding through the integration of rich meta-knowledge derived from multiple stages of the model development pipeline. This meta-knowledge is embedded within structured prompts, enabling the LLM to reason effectively over dataset characteristics, model properties, and optimization objectives. By combining informed contextual reasoning with iterative feedback, the framework not only identifies optimal hyperparameter settings but also exposes the underlying rationale guiding the optimization decisions. This section describes the construction of the meta-knowledge, the prompt design, and the overall optimization workflow.
3.1 Forecasting-Specific Meta-Knowledge
Meta-knowledge provides the contextual foundation for LLM-AutoOpt, enabling the LLM to reason over task characteristics, prior optimization behavior, and modeling constraints. Unlike conventional HPO methods that rely solely on objective evaluations, LLM-AutoOpt conditions optimization on structured information describing the forecasting problem and the optimization process itself.
Time-Series Meta-Features.
We extract a comprehensive set of meta-features to capture the statistical, temporal, and structural properties of the input time series. These include descriptive statistics (e.g., mean, variance, percentiles), distributional measures (skewness, kurtosis), and variability indicators (range, interquartile range, coefficient of variation). Temporal dependencies are quantified using autocorrelation and partial autocorrelation coefficients at multiple lags, while trend and seasonality strength are estimated via decomposition-based measures. Additional features describe data quality and complexity, including missing value ratios, outlier proportions, peak and trough counts, nonlinearity proxies, and stationarity diagnostics such as the Augmented Dickey–Fuller (ADF) statistic and p-value. Together, these meta-features provide a compact yet expressive representation of the forecasting task, allowing the LLM to adapt its recommendations to diverse temporal regimes.
Meta-Feature Statistical Summarization.
Beyond providing raw meta-features, we derive higher-level statistical summaries that aggregate these descriptors into interpretable indicators of the dataset behavior. Examples include summarizing autocorrelation coefficients across multiple lags to characterize overall temporal dependence, aggregating trend and stationarity statistics (e.g., ADF statistics and trend slopes) to describe global non-stationarity, and combining volatility- and noise-related measures (e.g., coefficient of variation, residual variance, and outlier ratio) to estimate the effective noise level of the series. These summarized representations are injected into the prompt in human-readable form, enabling the LLM to reason over dataset characteristics at a semantic level level rather than directly interpreting raw numerical statistics. This Abstraction facilitates better generalization in decision-making and significantly reduces hallucinated reasoning by constraining the LLM’s inference process to validate dataset-derived indicators.
Bayesian Optimization Initialization.
Bayesian Optimization (BO) is employed as an initial exploration mechanism to efficiently probe the hyperparameter space. The best-performing configuration—defined as the trial achieving the minimum observed validation loss. It is treated as task-specific meta-knowledge. Conditioning the LLM on this configuration provides a strong inductive bias, enabling it to refine an already high-quality solution rather than reasoning from an uninformed prior. This strategy improves sample efficiency, reduces convergence time, and mitigates the risk of stagnation in suboptimal regions of the search space.
Model Specification.
The model specification explicitly encodes the architecture and training setup of the candidate forecasting model, including layer types, hidden dimensions, activation functions, optimizers, learning rates, and loss functions. Providing this information as structured meta-knowledge grounds the LLM’s reasoning in the intended experimental design, reducing hallucination and ensuring that proposed hyperparameter updates remain consistent with the model’s functional constraints.
Search Space Constraints.
The hyperparameter search space is defined through explicit bounds and discrete choices for each parameter. These constraints serve both as a safety mechanism and as a reasoning aid, preventing invalid configurations (e.g., negative learning rates or unsupported batch sizes) while narrowing the space of plausible recommendations. Explicit constraints are essential for stable interaction between the LLM and the training pipeline. It is designed to act as a trust region centered around the current best-performing configuration.This formulation aims to reduce variance in the LLM’s recommendations while preserving controlled exploration of promising hyperparameter regions.
Few-Shot Task Context (Optional).
When available, a small number of representative input–output examples may be included to provide additional contextual grounding. These examples help the LLM reason about scale, variability, and noise characteristics of the data without relying on full dataset access.
Target Loss Criterion.
The optimization process is governed by a target loss threshold derived from the BO phase. Specifically, the target loss is defined as
$\text{Target\_Loss}=L_{\min}^{\text{BO}}-\varepsilon,$
where $L_{\min}^{\text{BO}}$ is the best loss observed during BO and $\varepsilon>0$ is a user-defined margin. This criterion serves as a stopping condition for the iterative LLM-guided optimization, ensuring that computational effort is focused on refining strong configurations while avoiding unnecessary evaluations.
3.2 Structured Prompt Design
The meta-knowledge is embedded within a structured framework, zero-shot prompt that conditions the LLM to generate hyperparameter updates. The prompt is designed to balance expressive reasoning with strict structural constraints. It enforces a single valid JSON output, prohibits duplicate hyperparameters, and requires adherence to predefined types and search space bounds. Internal reasoning fields encourage the LLM to justify each recommendation based on the provided meta-knowledge, improving transparency without exposing free-form text that could destabilize downstream processing. This design minimizes hallucinations, ensures consistency across iterations, and enables reliable integration of LLM outputs into the optimization loop.
3.3 LLM-AutoOpt Optimization Workflow
Figure 1 illustrates the overall optimization workflow, which consists of two tightly coupled phases.
Figure 1: Automated hyperparameter optimization guided by LLM reasoning.
Phase 1: Meta-Knowledge Construction.
Meta-features are first extracted from the input time-series data to characterize its temporal and statistical structure. Bayesian Optimization is then applied to explore the hyperparameter space of the candidate model. The best-performing trial, together with the extracted meta-features, model specification, and search space constraints, forms the initial meta-knowledge base.
Phase 2: LLM-Guided Iterative Refinement.
The meta-knowledge is injected into the structured prompt and provided to the LLM, which proposes refined hyperparameter configurations. The candidate model is trained using these recommendations, and the resulting performance metric (e.g., RMSE) is fed back into the prompt as updated context. This iterative loop continues until the target loss criterion is satisfied. The number of initial trials is dynamically increased whenever a hyperparameter configuration outperforms the current best RMSE identified by Bayesian optimization.
By integrating Bayesian initialization with LLM-based in-context reasoning, LLM-AutoOpt enables sample-efficient, interpretable, and forecasting-aware hyperparameter optimization. Unlike conventional HPO methods that operate on isolated trials, the framework allows the LLM to reason over historical optimization behavior and task-specific structure, leading to more informed and adaptive optimization trajectories.
4 Experimental Evaluation
This section evaluates the effectiveness of LLM-guided hyperparameter optimization, with a focus on both predictive performance and interpretability. In particular, we investigate how structured meta-knowledge influences optimization efficiency, model accuracy, and the transparency of the decision-making process. Our evaluation is guided by the following research questions:
  1. 1.
    RQ1 (Performance): How does the proposed LLM-AutoOpt framework compare to Bayesian Optimization in terms of forecasting accuracy (RMSE)?
  2. 2.
    RQ2 (Efficiency): How do training time and total optimization time differ between LLM-based HPO and BO-based baselines?
  3. 3.
    RQ3 (Optimization Dynamics): Does LLM-guided refinement converge to high-quality hyperparameter configurations with fewer evaluations than BO?
  4. 4.
    RQ4 (Role of Meta-Knowledge): How does incorporating forecasting-specific meta-knowledge affect the quality and stability of LLM-generated hyperparameter recommendations?
  5. 5.
    RQ5 (Interpretability): To what extent does LLM-AutoOpt improve interpretability by exposing reasoning signals and meta-knowledge usage compared to black-box optimization methods?
4.1 Experimental Setup
4.1.1 Dataset and Preprocessing
We conduct experiments on the Jena Climate multivariate time-series dataset, collected at the meteorological station of the Max Planck Institute for Biogeochemistry in Jena, Germany. The dataset contains high-frequency measurements recorded every 10 minutes between January 1, 2004 and December 31, 2020 and is widely used as a benchmark for multivariate forecasting tasks [24]. Each timestamp includes 14 meteorological variables such as air temperature, humidity, atmospheric pressure, and wind characteristics. To ensure computational tractability while preserving temporal structure, we extract a continuous six-month subset of the data. The CO2 variable is excluded due to extensive missing values. The remaining variables—air temperature (T), relative humidity (rh), atmospheric pressure (p), and wind speed (wv)—are selected for their relevance and complementary physical interpretations.
All variables are normalized using min–max scaling. The forecasting horizon is set to four time stamps, i.e., 40 minutes ahead. The data are split chronologically into training (70%) and testing (30%) sets to prevent information leakage across time.
4.1.2 Forecasting Model
Bidirectional LSTM (Bi-LSTM).
We select the Bidirectional Long Short-Term Memory (Bi-LSTM) network as the candidate forecasting model due to its widespread use and effectiveness in multivariate time-series modeling [25]. Bi-LSTMs process sequences in both forward and backward directions, enabling the capture of long-range temporal dependencies.
Optimized Hyperparameters.
We optimize a set of hyperparameters that strongly influence model performance and generalization, including batch size, hidden dimension, number of layers, dropout rate, learning rate, number of epochs, lag length, and optimizer choice. Table 1 summarizes these parameters and their effects. A randomly initialized baseline configuration (Table 2) is used for comparison.
Table 1: List of Optimized Hyperparameters
Hyperparameter Description & Impact
Batch size Number of samples per update; affects training stability and generalization
Hidden size Number of LSTM units; larger size captures complex patterns but may overfit
Number of layers Stacked LSTM layers; more layers capture long-term dependencies but may overfit
Dropout Fraction of neurons ignored during training; prevents overfitting, improves robustness
Learning rate (lr) Step size for weight updates; high lr speeds up training but may be unstable
Epochs Number of passes over data; more epochs improve learning but may overfit
Lag Number of previous time steps as input; captures temporal dependencies
Optimizer Method for updating weights (e.g., Adam, SGD); affects convergence and accuracy
Table 2: Baseline Hyperparameters for the Bi-LSTM Model
Hyperparameter Value
Hidden size 5
Number of layers 6
Dropout 0.5
Lag 3
Batch size 128
Epochs 10
Learning rate $1\times 10^{-9}$
Optimizer Adam
4.1.3 Large Language Model Backbone
The LLM backbone used for hyperparameter reasoning is Qwen2.5-72B, a large-scale language model developed by the Qwen team at Alibaba Group. The model is pretrained on multilingual and multimodal data and further aligned using high-quality instruction datasets. Its strong contextual reasoning capabilities make it well-suited for structured decision-making tasks such as hyperparameter optimization. Prior studies show that Qwen2.5-72B achieves low regret as the number of support tasks increases, outperforming alternative LLMs in optimization-oriented settings [22, 23]. All LLM outputs are validated against predefined constraints to ensure correctness and feasibility. To balance deterministic behavior and mitigate hallucinations, the decoding temperature parameter was set to 0.2.
4.1.4 Meta-Knowledge Prompt Configuration
The LLM prompt integrates structured meta-knowledge from all stages of the ML pipeline:
Data Meta-Knowledge.
Statistical, temporal, and structural meta-features describing the multivariate time series are provided (Table 3). These features summarize distributional properties, temporal dependencies, seasonality, stationarity, and data quality.
Table 3: Meta-features used to describe the input data
Meta-feature Description
count Number of observations in the column
missing Number of missing (NaN) values
mean Average value of the data
std Standard deviation of the data
min Minimum value
25% 25th percentile
50%_median Median (50th percentile)
75% 75th percentile
max Maximum value
skewness Measure of asymmetry of the distribution
kurtosis Measure of "tailedness" of the distribution
range Difference between max and min
iqr Interquartile range (75%-25%)
variance Variance of the data
acf_lag_1,3,6,12,24 Autocorrelation at specified lags
pacf_lag_1,3,6,12,24 Partial autocorrelation at specified lags
trend_strength Absolute slope of linear trend (regression)
adf_stat ADF test statistic for stationarity
adf_pvalue p-value of ADF test
num_peaks Number of local maxima
num_troughs Number of local minima
zero_ratio Fraction of zero values in series
outlier_ratio Fraction of outliers based on 1.5*IQR rule
coef_of_variation Relative variability (std/mean)
nonlinearity_proxy Std of first differences, indicates nonlinearity
trend_strength_decomp Strength of trend component from decomposition
seasonal_strength_decomp Strength of seasonal component from decomposition
residual_strength Variance of residual component from decomposition
Model Meta-Knowledge.
A detailed architectural description of the Bi-LSTM is included, specifying bidirectional processing, stacked layers, dropout usage, loss function, and optimization objective.
Optimization History.
The best-performing configurations identified by Bayesian Optimization are provided as historical trials (Table 4). These trials serve as task-specific meta-knowledge, enabling the LLM to refine strong initial solutions rather than exploring from scratch.
Table 4: Historical Bayesian Optimization trials used as contextual guidance for LLM-based hyperparameter recommendation
Trial ID Lag Hidden Size Layers Dropout LR Batch Epochs Optimizer
trial_3 57 56 1 0.50 9.84e-3 32 12 Adamax
trial_4 46 118 1 0.43 3.58e-3 64 19 AdamW
trial_1 36 39 1 0.27 8.44e-4 128 10 AdamW
Target Loss Criterion.
The optimization loop terminates when the RMSE reaches a target threshold defined as the best BO loss minus a margin $\varepsilon=0.09$, corresponding to a target RMSE of 1.00.
Prompt Constraints.
Strict output constraints enforce valid JSON formatting, prohibit repeated hyperparameters, and ensure compliance with search space boundaries. These constraints minimize hallucination and ensure that recommendations are directly executable.
4.1.5 Baselines and Method Variant
We evaluate LLM-AutoOpt against several baselines. First, we consider a baseline Bi-LSTM model with randomly initialized hyperparameters, denoted as Bi-LSTM-Base, using the configuration specified in Table 2. Second, we compare against a Bi-LSTM whose hyperparameters are selected via Bayesian Optimization, referred to as Bi-LSTM-BO. Finally, we assess an ablated variant of our method that excludes meta-knowledge from the prompt, denoted as LLM-AutoOpt-NoMeta, to isolate the impact of meta-knowledge on optimization performance.
4.1.6 Evaluation Metrics
We evaluate performance using three complementary metrics:
Forecasting Accuracy (RMSE).
Root Mean Square Error (RMSE) is used to assess predictive accuracy. It penalizes large errors and is well suited for continuous forecasting tasks. RMSE serves both as the evaluation metric and as feedback for the LLM-guided optimization loop.
Training Time.
Training time measures the computational cost of fitting the model given a fixed hyperparameter configuration. This metric captures the downstream cost induced by different optimization strategies.
Optimization Time.
Optimization time measures the total wall-clock time required to complete the hyperparameter optimization process, including both BO and LLM-guided iterations.
4.2 Results
This section presents a quantitative and qualitative evaluation of the proposed LLM-AutoOpt framework. Results are organized around the research questions introduced in Section 4, with emphasis on predictive performance, optimization efficiency, convergence behavior, and interpretability.
4.2.1 RQ1: Predictive Performance
Figure 2 and Table 5 compare forecasting accuracy across all optimization strategies using RMSE as the primary metric. LLM-AutoOpt consistently achieves the lowest RMSE among all baselines, including Bayesian Optimization (Bi-LSTM-BO), an LLM without meta-knowledge (LLM-AutoOpt-NoMeta), and the manually initialized baseline model (Bi-LSTM-Base). LLM-AutoOpt exhibits faster convergence and stabilizes at a significantly lower error level throughout training. In contrast, Bi-LSTM-BO converges more slowly and fails to reach the predefined target RMSE, remaining approximately $0.09$ above the threshold. The LLM-AutoOpt-NoMeta improves over the baseline but demonstrates higher variance and inferior final accuracy, highlighting the limitations of unguided LLM reasoning.
Figure 2: Validation RMSE convergence across optimization strategies
Table 5: Comparison of Mean RMSE, Average Training Time, and Optimization Time for Different Methods (with Standard Deviations)
Method Mean RMSE RMSE Std Training Time (s) Optimization Time (s)
Bi-LSTM-Base 5.42 0.31 31 -
Bi-LSTM-BO (Top 3) 1.19 0.03 30 82
LLM-AutoOpt-NoMeta(Top 3) 1.90 0.05 549 1065
LLM-AutoOpt(Top 3) 1.11 0.04 46 177
Beyond aggregate error metrics, Figure 3 provides a qualitative comparison of the relative humidity (RH %) forecasts produced by different models. The forecasts generated by LLM-AutoOpt more closely track the ground-truth signal, capturing both local fluctuations and longer-term temporal trends. Competing methods exhibit noticeable smoothing effects or delayed responses to rapid changes, indicative of suboptimal hyperparameter configurations.
Figure 3: Relative humidity (RH %) forecast comparison across different models
Taken together, these quantitative and qualitative results demonstrate that meta-knowledge–guided LLM optimization not only reduces overall forecasting error but also improves the fidelity of time-series predictions, particularly in capturing dynamic temporal patterns.
4.2.2 RQ2: Computational Efficiency
Table 5 reports both training time and total optimization time. Bayesian Optimization remains the most computationally efficient approach, requiring the least optimization time and minimal per-trial training cost. In contrast, LLM-based methods incur additional overhead due to prompt construction, reasoning, and iterative feedback. Notably, LLM-AutoOpt is significantly more efficient than the LLM-AutoOpt-NoMeta variant, reducing optimization time by more than $80\%$ while achieving substantially better accuracy. This indicates that meta-knowledge not only improves solution quality but also constrains the search space effectively, reducing unnecessary exploration. While LLM-AutoOpt does not outperform Bi-LSTM-BO in raw computational speed, it offers a different trade-off: improved accuracy and interpretability at the cost of moderate additional overhead.
4.2.3 RQ3: Optimization Dynamics and Convergence Behavior
The RMSE trajectories in Figure 2 reveal distinct optimization dynamics. LLM-AutoOpt exhibits fast early-stage improvement followed by stable convergence, suggesting that the LLM leverages meta-knowledge to make informed, high-impact adjustments early in the optimization process. In contrast, the LLM-AutoOpt-NoMeta shows erratic convergence behavior, often relying on brute-force strategies such as increasing the number of epochs to reduce RMSE. Bi-LSTM-BO follows a smoother but slower trajectory and appears prone to local stagnation near suboptimal regions. These observations suggest that LLM-guided refinement, when grounded in meta-knowledge, can escape limitations of purely surrogate-based optimization and converge with fewer ineffective trials.
4.2.4 RQ4: Impact of Meta-Knowledge on Optimization Quality
Table 6 highlights qualitative differences in the hyperparameter configurations produced by each method. LLM-AutoOpt often proposes configurations that resemble strong BO solutions (e.g., similar batch sizes or layer counts) while introducing targeted refinements such as reducing dropout, adjusting lag length, or lowering epoch counts to mitigate overfitting.
Table 6: Hyperparameter configurations and RMSE results across optimization trials with method comparisons
Trial Method Lag Hidden Size Layers Dropout LR Batch Epochs Optimizer RMSE
1 Bi-LSTM-BO 36 39 1 0.27 8.44e-4 128 10 AdamW 1.42
LLM-AutoOpt-NoMeta 24 50 3 0.20 1.00e-3 64 100 Adam 2.10
LLM-AutoOpt 12 64 2 0.15 1.00e-3 64 40 Adam 1.15
2 Bi-LSTM-BO 57 56 1 0.50 9.84e-3 32 12 Adamax 1.09
LLM-AutoOpt-NoMeta 24 100 5 0.30 5.00e-3 128 200 RMSprop 1.95
LLM-AutoOpt 16 48 2 0.10 1.00e-3 64 30 Adam 1.18
3 Bi-LSTM-BO 46 118 1 0.43 3.58e-3 64 19 AdamW 1.09
LLM-AutoOpt-NoMeta 24 100 5 0.30 5.00e-3 128 200 RMSprop 1.66
LLM-AutoOpt 12 64 1 0.10 1.00e-3 64 40 Adam 1.01
By contrast, the LLM-AutoOpt-NoMeta frequently selects extreme configurations, particularly large epoch counts, as a generic mechanism for RMSE reduction. This behavior indicates a lack of structural understanding of the task and results in inefficient or unstable optimization. These findings confirm that historical trials and model-aware context function as actionable meta-knowledge, enabling the LLM to prioritize meaningful adjustments rather than arbitrary exploration.
4.2.5 RQ5: Interpretability of LLM-Guided Optimization
Figure 4 contrasts reasoning traces generated with and without meta-knowledge injection. The LLM-AutoOpt-NoMeta produces abstract and generic explanations, typically referencing prior RMSE values without explicit justification tied to model architecture or data properties. In contrast, LLM-AutoOpt provides grounded and task-aware reasoning, explicitly linking hyperparameter changes to architectural considerations (e.g., reducing dropout to alleviate underfitting) or temporal characteristics of the data. This transparency enables practitioners to audit, trust, and potentially reuse optimization insights—an advantage absent in black-box methods such as BO. Overall, meta-informed reasoning improves not only optimization outcomes but also the interpretability and explanatory power of the optimization process.
Figure 4: Comparison of reasoning outputs for hyperparameter recommendation with and without meta-knowledge injection.
4.3 Discussion
The experimental results demonstrate that incorporating meta-knowledge into LLM-based hyperparameter optimization yields consistent improvements in forecasting accuracy and optimization behavior. Across all evaluated settings, LLM-AutoOpt achieves lower RMSE and faster convergence compared to manually initialized models, Bayesian Optimization, and LLM-based optimization without meta-knowledge. These gains are reflected not only in quantitative metrics but also in the qualitative forecasting results, where meta-informed configurations produce smoother and more accurate predictions that better capture temporal dynamics.
From a methodological perspective, the results indicate that meta-knowledge plays a dual role. First, it improves optimization effectiveness by providing the LLM with structured priors over the data distribution, model architecture, and historical optimization outcomes. This reduces unproductive exploration and allows the LLM to focus on meaningful hyperparameter adjustments, such as tuning lag length, regularization strength, or network depth. Second, meta-knowledge stabilizes the reasoning process itself. In contrast to the LLM-AutoOpt-NoMeta setting, where the LLM often relies on generic heuristics (e.g., increasing epochs or model size), LLM-AutoOpt generates decisions that are explicitly grounded in dataset characteristics and model constraints.
Interpretability emerges as a key advantage of the proposed framework. The reasoning traces produced by LLM-AutoOpt reveal how specific meta-features and architectural considerations influence hyperparameter choices, enabling a transparent link between data properties and optimization decisions. This stands in contrast to Bayesian Optimization, which, while computationally efficient, offers limited insight into why particular configurations are preferred. As such, LLM-AutoOpt provides a more explainable alternative for practitioners seeking both performance and interpretability.
However, these benefits come at the cost of computational efficiency. BO remains substantially faster in terms of both optimization and training time, highlighting a trade-off between speed and interpretability. This suggests that LLM-AutoOpt is best suited for scenarios where accuracy, robustness, and explainability are prioritized over raw optimization speed. Additionally, the framework’s performance depends on the quality of prompt design and the relevance of extracted meta-features, indicating that careful engineering is required to fully realize its benefits.
Finally, while the results are encouraging, they are based on a single multivariate forecasting benchmark. Further evaluation across diverse datasets, architectures, and forecasting horizons is necessary to assess the generality and scalability of the approach.
5 Conclusion
This work presented LLM-AutoOpt, a meta-knowledge–guided hyperparameter optimization framework that leverages large language model reasoning for multivariate time-series forecasting. By integrating dataset meta-features, model descriptions, and historical optimization results into a structured prompting strategy, the framework enables more accurate, stable, and interpretable hyperparameter tuning. Empirical results show that LLM-AutoOpt consistently outperforms traditional Bayesian Optimization and LLM-based baselines without meta-knowledge in terms of predictive accuracy, while providing transparent and data-aware optimization reasoning. Although the approach does not improve computational efficiency, it offers a principled and explainable alternative for high-stakes forecasting tasks.
Future work will explore cross-dataset generalization, automated meta-feature selection, and extensions to additional forecasting domains and deep sequence models, further advancing the role of meta-knowledge–augmented LLMs in hyperparameter optimization.